# Turbo-GS: Accelerating 3D Gaussian Fitting for High-Quality Radiance Fields

### v1

@@ 5.3 Results and Analysis
For quantitative analysis, we followed evaluation methodology proposed in 3DGS [24]. We observe that 3DGS takes on an average of $\sim$30 minutes for MipNeRF-360 scenes in Table 1. Other 3DGS-based methods like EAGLES and Mini-Splatting are faster than 3DGS but take about $\sim$20 minutes. Taming 3DGS is fast, but it still takes $\sim$16 minutes on these challenging scenes. In contrast, Turbo-GS takes only $\sim$5 minutes which is 3$\times$ faster than Taming 3DGS and 5.6$\times$ faster than 3DGS methods. In addition, we maintain similar or better quality than the other methods. The average training time of Turbo-GS for Deep Blending is $\sim$3 minutes.
@@
Table 2: Quantitative comparison on 3DGS-based methods for MipNeRF360-4K [1]. Turbo-GS (Ours) is not only fast but outperforms baseline methods in all the metrics.
@@
3DGS [24] 26.75 0.797 0.410 113 m
Taming 3DGS [30] 26.66 0.795 0.357 28 m
Scaffold-GS [29] 26.84 0.794 0.359 143 m
@@ 5.4 Ablation Study
We present ablation studies on key design choices in our method. More ablations are in the supplementary material. All the ablation are conducted on the bicycle scene from MipNeRF360 [1].
@@
We show how cues from color gradients support position gradients during densification We present two strategies: (1) position only and (2) position $+$ color($p$), where “$p$” denotes the probability of enabling color gradients, as shown in Table 3. While the position-only design achieves reasonable PSNR, it significantly affects image structure, as indicated by a low LPIPS score. In contrast, introducing color gradients improves LPIPS, validating the effectiveness of the proposed densification strategy. We observe that the color-gradient’s frequency has little impact on quality.
Table 3: Ablation on importance of color gradients in densification strategy for Bicycle [1]. $p$ denotes the probability of enabling color gradient.
@@
Table 4: Ablation on Dilated Rendering for Bicycle [1].
@@ 6 Conclusion and Limitations
In this work, we present Turbo-GS, an efficient method that accelerates 3DGS model fitting by 5×5$\times$5 × compared to previous methods. We propose two major design choices in the current optimization framework. First, we design an aggressive densification strategy that achieves faster optimization and requires fewer steps. Our densification strategy performs better than other methods in textured regions such as grass, floors, etc. Secondly, we introduce dilated rendering, which renders only a subset of the pixels rather than the entire image, allowing accelerated training on images exceeding 4K resolution while maintaining high quality. Our approach outperforms all existing methods in training time for standard 1K scenes while matching or surpassing their rendering quality. We also exceed baseline methods on existing 4K datasets. This paper aims at a learning-free design, thus lacking exploration on how to integrate with a geometry foundation model. With many such models gradually changing the entire community, we will consider how to benefit from a large model in the future to achieve further acceleration.
@@ 7.2 Convergence-aware Budget Control
Figure 8: Convergence. We show “Number of primitives vs Step” and “PSNR vs Step” plots for scenes in MipNeRF-360 [1] dataset for with and without budget control in the optimization process. The proposed budgeting strategy prevents the number of primitives from increasing uncontrollably, while maintaining the overall quality. This is evident by the comparable PSNR plots, which demonstrate that the strategy maintains the balance between computational efficiency and visual fidelity.
Figure 9: Impact of Dilated Rendering on time performance. We observe that dilated rendering significantly reduces the computational time required for both the (a) forward and (b) backward passes during the optimization process, compared to the without-dilated rendering approach. This highlights the efficiency of dilated rendering in accelerating the overall training process. The above results are shown for Bicycle from MipNeRF [1]. This enables Turbo-GS to achieve faster fitting for $\geq$4K scenes.
@@
0:   $steps$: Total steps
0:   $window\_size$: Window for trend analysis
@@
13:      else if $rate<-0.05$ then
@@ 8 More Experiments and Results
Here we list the error metrics used in our evaluation in Sec.4 across all considered methods and scenes, as shown in Tab 58. drjohnson-playroom [19] belongs to the deep blending dataset; train-truck come from the Tanks and Temple [27] dataset; bicycle-boonsai are from MipNeRF360 [1].
@@
Table 5: PSNR comparison across DeepBlending [19] (“drjohnson”, “playroom”), Tanks & Temples [27] (“train”, “truck”) and MipNeRF-360 [1] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [24] 29.49 30.02 22.11 25.45 25.23 27.38 26.59 21.44 22.49 29.08 31.09 31.48 32.31 Taming-3DGS [30] 29.39 30.04 22.09 25.46 25.22 27.35 26.62 21.50 22.59 29.07 30.98 31.62 32.22 Taming-3DGS* [30] 29.36 29.94 21.77 25.29 25.13 27.24 26.46 21.43 22.53 28.95 31.09 31.18 32.03 Mini-Splatting [12] 29.35 30.46 21.46 25.06 25.22 26.85 27.22 21.60 22.68 28.60 31.31 31.36 31.21 EAGLES [16] 29.35 30.19 21.36 25.00 24.97 26.87 26.61 21.31 22.62 28.30 30.52 31.38 31.23 ScaffoldGS [29] 29.67 30.90 22.53 25.86 25.21 27.51 26.56 21.43 23.12 29.45 31.61 31.98 32.53 Mip-Splatting [47] 28.75 29.98 22.06 25.74 25.56 27.69 26.91 21.70 22.35 29.13 31.59 31.54 32.27 Ours 29.69 31.12 21.31 25.67 25.01 27.43 26.41 21.12 22.75 29.22 30.96 31.23 32.29
@@
Table 6: SSIM comparison across DeepBlending [19] (“drjohnson”, “playroom”), Tanks & Temples [27] (“train”, “truck”) and MipNeRF-360 [1] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [24] 0.903 0.902 0.818 0.881 0.765 0.864 0.770 0.602 0.633 0.907 0.925 0.918 0.940 Taming-3DGS [30] 0.902 0.904 0.818 0.881 0.765 0.863 0.770 0.601 0.634 0.906 0.925 0.919 0.939 Taming-3DGS* [30] 0.903 0.901 0.812 0.878 0.751 0.859 0.764 0.596 0.630 0.905 0.923 0.915 0.939 Mini-Splatting [12] 0.903 0.912 0.798 0.874 0.773 0.848 0.806 0.626 0.654 0.905 0.926 0.922 0.939 EAGLES [16] 0.906 0.908 0.796 0.872 0.757 0.844 0.770 0.589 0.637 0.897 0.920 0.917 0.934 ScaffoldGS [29] 0.907 0.912 0.822 0.886 0.760 0.863 0.766 0.594 0.643 0.911 0.927 0.924 0.944 Mip-Splatting [47] 0.898 0.908 0.827 0.893 0.793 0.878 0.791 0.640 0.639 0.913 0.930 0.925 0.944 Ours 0.905 0.915 0.795 0.887 0.754 0.863 0.761 0.593 0.627 0.912 0.927 0.921 0.945
@@
Table 7: LPIPS comparison across DeepBlending [19] (“drjohnson”, “playroom”), Tanks & Temples [27] (“train”, “truck”) and MipNeRF-360 [1] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [24] 0.238 0.243 0.198 0.143 0.211 0.108 0.217 0.339 0.329 0.201 0.127 0.220 0.205 Taming-3DGS [30] 0.239 0.243 0.200 0.144 0.210 0.109 0.217 0.341 0.328 0.202 0.127 0.219 0.206 Taming-3DGS* [30] 0.244 0.253 0.206 0.147 0.236 0.116 0.229 0.346 0.341 0.204 0.130 0.229 0.208 Mini-Splatting 0.256 0.249 0.245 0.160 0.225 0.150 0.199 0.327 0.313 0.198 0.129 0.211 0.200 EAGLES [16] 0.244 0.253 0.240 0.166 0.232 0.146 0.229 0.361 0.338 0.217 0.138 0.226 0.218 ScaffoldGS [29] 0.252 0.253 0.206 0.142 0.227 0.118 0.236 0.347 0.319 0.200 0.127 0.210 0.203 Mip-Splatting [47] 0.243 0.235 0.189 0.123 0.167 0.094 0.188 0.274 0.274 0.187 0.119 0.202 0.188 Ours 0.249 0.230 0.219 0.133 0.223 0.113 0.228 0.317 0.302 0.190 0.124 0.204 0.191
@@
Table 8: Training time (in minutes) comparison across DeepBlending [19] (“drjohnson”, “playroom”), Tanks & Temples [27] (“train”, “truck”) and MipNeRF-360 [1] scenes.

### v2

@@ 5.3 Results and Analysis
[…]
@@
Table 2: Quantitative comparison on 3DGS-based methods for MipNeRF360-4K [mip360]. Turbo-GS (Ours) is not only fast but outperforms baseline methods in all the metrics.
@@
3DGS [3dgs] 26.75 0.797 0.410 113 m
Taming 3DGS [taming3dgs] 26.66 0.795 0.357 28 m
Scaffold-GS [scaffoldgs] 26.84 0.794 0.359 143 m
@@ 5.4 Ablation Study
We present ablation studies on key design choices in our method. More ablations are in the supplementary material. All the ablation are conducted on the bicycle scene from MipNeRF360 [mip360].
@@
We show how cues from color gradients support position gradients during densification We present two strategies: (1) position only and (2) position ++ color($p$), where “pp” denotes the probability of enabling color gradients, as shown in Table 3. While the position-only design achieves reasonable PSNR, it significantly affects image structure, as indicated by a low LPIPS score. In contrast, introducing color gradients improves LPIPS, validating the effectiveness of the proposed densification strategy. We observe that the color-gradient’s frequency has little impact on quality.
Table 3: Ablation on importance of color gradients in densification strategy for Bicycle [mip360]. $p$ denotes the probability of enabling color gradient.
@@
Table 4: Ablation on Dilated Rendering for Bicycle [mip360].
@@ 6 Conclusion and Limitations
In this work, we present Turbo-GS, an efficient method that accelerates 3DGS model fitting by $5\times$ compared to previous methods. We propose two major design choices in the current optimization framework. First, we design an aggressive densification strategy that achieves faster optimization and requires fewer steps. Our densification strategy performs better than other methods in textured regions such as grass, floors, etc. Secondly, we introduce dilated rendering, which renders only a subset of the pixels rather than the entire image, allowing accelerated training on images exceeding 4K resolution while maintaining high quality. Our approach outperforms all existing methods in training time for standard 1K scenes while matching or surpassing their rendering quality. We also exceed baseline methods on existing 4K datasets. This paper aims at a learning-free design, thus lacking exploration on how to integrate with a geometry foundation model. With many such models gradually changing the entire community, we will consider how to benefit from a large model in the future to achieve further acceleration.
@@ 7.2 Convergence-aware Budget Control
Figure 8: Convergence. We show “Number of primitives vs Step” and “PSNR vs Step” plots for scenes in MipNeRF-360 [mip360] dataset for with and without budget control in the optimization process. The proposed budgeting strategy prevents the number of primitives from increasing uncontrollably, while maintaining the overall quality. This is evident by the comparable PSNR plots, which demonstrate that the strategy maintains the balance between computational efficiency and visual fidelity.
Figure 9: Impact of Dilated Rendering on time performance. We observe that dilated rendering significantly reduces the computational time required for both the (a) forward and (b) backward passes during the optimization process, compared to the without-dilated rendering approach. This highlights the efficiency of dilated rendering in accelerating the overall training process. The above results are shown for Bicycle from MipNeRF [mip360]. This enables Turbo-GS to achieve faster fitting for $\geq$4K scenes.
@@
0:  stepssteps: Total steps
0:  window_sizewindow\_size: Window for trend analysis
@@
13:   else if rate<0.05rate<-0.05 then
@@ 8 More Experiments and Results
Here we list the error metrics used in our evaluation in Sec.4 across all considered methods and scenes, as shown in Tab 58. drjohnson-playroom [hedman2018deep] belongs to the deep blending dataset; train-truck come from the Tanks and Temple [knapitsch2017tanks] dataset; bicycle-boonsai are from MipNeRF360 [mip360].
@@
Table 5: PSNR comparison across DeepBlending [hedman2018deep] (“drjohnson”, “playroom”), Tanks & Temples [knapitsch2017tanks] (“train”, “truck”) and MipNeRF-360 [mip360] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [3dgs] 29.49 30.02 22.11 25.45 25.23 27.38 26.59 21.44 22.49 29.08 31.09 31.48 32.31 Taming-3DGS [taming3dgs] 29.39 30.04 22.09 25.46 25.22 27.35 26.62 21.50 22.59 29.07 30.98 31.62 32.22 Taming-3DGS* [taming3dgs] 29.36 29.94 21.77 25.29 25.13 27.24 26.46 21.43 22.53 28.95 31.09 31.18 32.03 Mini-Splatting [fang2024mini] 29.35 30.46 21.46 25.06 25.22 26.85 27.22 21.60 22.68 28.60 31.31 31.36 31.21 EAGLES [Girish2023EAGLESEA] 29.35 30.19 21.36 25.00 24.97 26.87 26.61 21.31 22.62 28.30 30.52 31.38 31.23 ScaffoldGS [scaffoldgs] 29.67 30.90 22.53 25.86 25.21 27.51 26.56 21.43 23.12 29.45 31.61 31.98 32.53 Mip-Splatting [Yu2023MipSplattingA3] 28.75 29.98 22.06 25.74 25.56 27.69 26.91 21.70 22.35 29.13 31.59 31.54 32.27 Ours 29.69 31.12 21.31 25.67 25.01 27.43 26.41 21.12 22.75 29.22 30.96 31.23 32.29
@@
Table 6: SSIM comparison across DeepBlending [hedman2018deep] (“drjohnson”, “playroom”), Tanks & Temples [knapitsch2017tanks] (“train”, “truck”) and MipNeRF-360 [mip360] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [3dgs] 0.903 0.902 0.818 0.881 0.765 0.864 0.770 0.602 0.633 0.907 0.925 0.918 0.940 Taming-3DGS [taming3dgs] 0.902 0.904 0.818 0.881 0.765 0.863 0.770 0.601 0.634 0.906 0.925 0.919 0.939 Taming-3DGS* [taming3dgs] 0.903 0.901 0.812 0.878 0.751 0.859 0.764 0.596 0.630 0.905 0.923 0.915 0.939 Mini-Splatting [fang2024mini] 0.903 0.912 0.798 0.874 0.773 0.848 0.806 0.626 0.654 0.905 0.926 0.922 0.939 EAGLES [Girish2023EAGLESEA] 0.906 0.908 0.796 0.872 0.757 0.844 0.770 0.589 0.637 0.897 0.920 0.917 0.934 ScaffoldGS [scaffoldgs] 0.907 0.912 0.822 0.886 0.760 0.863 0.766 0.594 0.643 0.911 0.927 0.924 0.944 Mip-Splatting [Yu2023MipSplattingA3] 0.898 0.908 0.827 0.893 0.793 0.878 0.791 0.640 0.639 0.913 0.930 0.925 0.944 Ours 0.905 0.915 0.795 0.887 0.754 0.863 0.761 0.593 0.627 0.912 0.927 0.921 0.945
@@
Table 7: LPIPS comparison across DeepBlending [hedman2018deep] (“drjohnson”, “playroom”), Tanks & Temples [knapitsch2017tanks] (“train”, “truck”) and MipNeRF-360 [mip360] scenes.
@@
Method drjohnson playroom train truck bicycle garden stump flowers treehill counter kitchen room bonsai 3DGS [3dgs] 0.238 0.243 0.198 0.143 0.211 0.108 0.217 0.339 0.329 0.201 0.127 0.220 0.205 Taming-3DGS [taming3dgs] 0.239 0.243 0.200 0.144 0.210 0.109 0.217 0.341 0.328 0.202 0.127 0.219 0.206 Taming-3DGS* [taming3dgs] 0.244 0.253 0.206 0.147 0.236 0.116 0.229 0.346 0.341 0.204 0.130 0.229 0.208 Mini-Splatting 0.256 0.249 0.245 0.160 0.225 0.150 0.199 0.327 0.313 0.198 0.129 0.211 0.200 EAGLES [Girish2023EAGLESEA] 0.244 0.253 0.240 0.166 0.232 0.146 0.229 0.361 0.338 0.217 0.138 0.226 0.218 ScaffoldGS [scaffoldgs] 0.252 0.253 0.206 0.142 0.227 0.118 0.236 0.347 0.319 0.200 0.127 0.210 0.203 Mip-Splatting [Yu2023MipSplattingA3] 0.243 0.235 0.189 0.123 0.167 0.094 0.188 0.274 0.274 0.187 0.119 0.202 0.188 Ours 0.249 0.230 0.219 0.133 0.223 0.113 0.228 0.317 0.302 0.190 0.124 0.204 0.191
@@
Table 8: Training time (in minutes) comparison across DeepBlending [hedman2018deep] (“drjohnson”, “playroom”), Tanks & Temples [knapitsch2017tanks] (“train”, “truck”) and MipNeRF-360 [mip360] scenes.